OpenStudy (anonymous):

Will Medal who awnsers correctly. 1. An item is priced at $13.16. If the sales tax is 4%, what does the item cost including sales tax? $13.69 $0.53 $18.42 $26.85

OpenStudy (gschibby):

Price with sales tax = price * 1,04

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@iGreen can you helps me?

OpenStudy (anonymous):

@amistre64 @iGreen

OpenStudy (anonymous):

@RyleeC

OpenStudy (anonymous):

i think it is a

OpenStudy (anonymous):

you all get a fan

OpenStudy (anonymous):

Heres how to do it...

OpenStudy (anonymous):

ik now :p

OpenStudy (anonymous):

okay nvmd

OpenStudy (igreen):

You have to remove the 4% sales tax from 13.16.

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

.o.

OpenStudy (anonymous):

Percent to number just move decimal by 2 to the left

OpenStudy (anonymous):

2. Sandy works at a clothing store. She makes $7 per hour plus earns 10% commission on her sales. She worked 71 hours over the last two weeks and had a total of $1,986 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? $497 $696 $199 $248

OpenStudy (anonymous):

because 13.16 +4%=13.6864 then you round to the NEAREST HUNDREDTH PLACE AND THEN IT 13.69

OpenStudy (anonymous):

.o.

OpenStudy (anonymous):

71*7=497 but

OpenStudy (anonymous):

;p

OpenStudy (anonymous):

.10*1986=198.6

OpenStudy (anonymous):

so the anser it a?

OpenStudy (anonymous):

No its B you have to add the two

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

Gj ryleec

OpenStudy (anonymous):

you are awesome at this

OpenStudy (anonymous):

Thanks, as long as you understand how I got the answer...

OpenStudy (anonymous):

;p

OpenStudy (anonymous):

this one is harder 3. A bakery sold 105 cupcakes in one day. The head baker predicted he would sell 85 cupcakes that day. What was the percent error of the baker's prediction? 23.53% 20% 80.95% 19.05%

OpenStudy (anonymous):

you have to get the %

OpenStudy (anonymous):

i think it is c

OpenStudy (anonymous):

wait no 20

OpenStudy (anonymous):

I BELIEVE ITS A

OpenStudy (anonymous):

i subtracted 105-85 and got 20

OpenStudy (gschibby):

No, c is correct ;) 85/105=0,8095=80,95% ;)

OpenStudy (anonymous):

let m,e see i will pick ant tell you

OpenStudy (gschibby):

You have to find out how many percent 85 is out of 105 ;)

OpenStudy (anonymous):

nope we were all wrong it was D xD

OpenStudy (anonymous):

I UNDERSTAND YOU SUBTRACTED BUT TO GET A PERCENT YOU DIVIDE

OpenStudy (anonymous):

What was the answer

OpenStudy (anonymous):

it was d

OpenStudy (anonymous):

The salaries of three employees are listed below. Each employee was recently given a raise. Employee Salaries Employee Salary [before raise] Raise Jane $41,100 4% John $43,300 4.5% Jack $39,900 3% Which amount is closest to the total of all three raises? $4,790 $14,295 $4,723 $10,566

OpenStudy (gschibby):

Oops, my bad, it's the other way around 105/85=1,2352 wich is 0,2352=23,52% above the estimate ;)

OpenStudy (anonymous):

it was a

OpenStudy (anonymous):

and i got i right :D ty ryleec